What is the angular velocity of an object traveling in a circle of radius 0.75 with a linear speed of 3.5m/s?

1
Expert's answer
2021-12-07T20:11:05-0500

The angular velocity "\omega" is "\omega = \dfrac{V}{R} = \dfrac{3.5\,\mathrm{m/s}}{0.75\,\mathrm{m}} \approx 4.67\,\mathrm{rad/s}."

Why is this formula correct? The angular velocity "\omega" is "\omega = \dfrac{2\pi}{T} = \dfrac{2\pi}{2\pi R/V} = \dfrac{V}{R}."